OpenStudy (anonymous):

derivative of: (4e^t-3t^2)

OpenStudy (anonymous):

Im not sure how to do the product rule for this

OpenStudy (across):

You're given:\[\frac{d}{dt}[4e^{t}-3t^{2}].\]Notice that this can be re-arranged as\[\frac{d}{dt}[4e^{t}]-\frac{d}{dt}[3t^{2}].\]

OpenStudy (anonymous):

yes i got that far

OpenStudy (across):

Do you know what the derivative of\[4e^{t}\]is?

OpenStudy (anonymous):

isnt it just 4e^t because the der. of e^t is e^t right?

OpenStudy (across):

Correct!

OpenStudy (across):

Now, what's the derivative of\[3t^2?\]

OpenStudy (anonymous):

before you get too far, is this \[4e^t-3t^2\] or \[4e^{t-3t^2}\] because it will surely make a difference

OpenStudy (anonymous):

the first one, satellite

OpenStudy (anonymous):

"product rule" suggested it might be the second one. of course i could be totally wrong, in which case go right along with across and i will be quiet

OpenStudy (anonymous):

ok then i shut up

OpenStudy (anonymous):

if you can help me then please go right ahead

OpenStudy (anonymous):

you have it. no product rule, just take the derivative term by term